Ok so i need help on finding the volume of a:
-rectangular prism,(length=12 in.,width=8 in.,hieght=7 in.)
-pyramid,(L=10 in.,W=9 in., H=12 in.)
-cone,(radius=7 in., Height=14 in.)
PLEASE HELP ME WITH THIS QUESTION AND IF YOU KAN, KAN U GET THIS BY ME TODAY B4 4 o' clock
THANK YOU :)

Answer by user_dude2008(1825)

#1

V = L*W*H = 12*8*7 = 96*7 = 672

Volume: 672 cubic inches
#2

V = (1/3)*(L*W)*H = (1/3)*(10*9)*12 = (1/3)*(90)*12 = 30*12 = 360

Volume: 360 cubic inches


#3

V = (1/3)pi*r^2*H = (1/3)(3.14)(7^2)(14) = 718.013

Volume: about 718.013 cubic inches